# Direct reconstruction of anisotropic self-adjoint inclusions in the Calderón problem

### v1

@@ Abstract.
We extend the monotonicity method for direct exact reconstruction of inclusions in the partial data Calderón problem, to the case of anisotropic conductivities in any spatial dimension $d\geq 2$. Specifically, from a local Neumann-to-Dirichlet map, we give reconstruction methods of inclusions based on unknown anisotropic self-adjoint perturbations to a known anisotropic conductivity coefficient. A main assumption is a definiteness condition for the perturbations near the outer inclusion boundaries. This additionally provides new insights into the non-uniqueness issues of the anisotropic Calderón problem.
@@ 1. Introduction
produces the same local Neumann-to-Dirichlet (ND) map as the coefficient $A$. For $\Gamma=\partial\Omega$ and with spatial dimension $d=2$, the obstructions to uniqueness are characterized by (1.1) for real-valued symmetric $L^{\infty}$-conductivities [4, 37]. In spatial dimension $d\geq 3$, the full extent of the non-uniqueness issues (beyond real-analytic coefficients) remains one of the most important unanswered questions in the inverse problems community [32, 39]. In a fractional Calderón problem, using the stronger non-local properties, the obstructions to uniqueness are characterized for $d\geq 2$ [9]. Despite knowledge on the possible obstructions, the non-uniqueness has prevented development of practically useful reconstruction methods. Unique recovery is only guaranteed with very strong restrictions on the considered class of coefficients [2, 3, 10].
@@
from knowledge of $A_{0}$ and the local ND map $\Lambda(A_{D})$ associated with the unknown conductivity $A_{D}$. Here $D^{\bullet}$ is the smallest closed set with connected complement such that $D\subseteq D^{\bullet}$. We give direct reconstruction methods (Theorems 3.6 and 3.7) based on the monotonicity of ND maps. The main assumptions are self-adjointness of the coefficients, that $A_{0}$ satisfies a unique continuation principle, and that $A_{D}-A_{0}$ satisfies a local uniform definiteness condition inside $D$ near $\partial D^{\bullet}$; see Assumption 3.4 for the precise statement. In particular, inside $D$ we do not assume more regularity than $L^{\infty}$ for $A_{D}$, and $A_{D}-A_{0}$ can even be indefinite at a positive distance away from $\partial D^{\bullet}$.
As an indirect consequence: if such a perturbation $A_{D}-A_{0}$, to a given “background conductivity” $A_{0}$, is known to satisfy the definiteness condition near $\partial D^{\bullet}$, this is enough to ensure that all coefficient-transformations (1.1), compatible with this condition, satisfy
$\mathop{\textup{supp}}(\widetilde{A}_{D}-A_{0})^{\bullet}=D^{\bullet}.$
Moreover, the inclusion detection problem is also unaffected by any other potential forms of non-uniqueness that could arise from the anisotropic Calderón problem when $d\geq 3$. Hence, for inclusion detection, practical reconstruction is sensible in a very high generality compared to reconstructing an anisotropic conductivity.
To give some intuition on why this is the case, let
@@
be the transformed inclusion under $\Phi$. If $\widetilde{D}^{\bullet}\neq D^{\bullet}$ then $\Phi$ will fail to be the identity in an open set intersecting $\partial D^{\bullet}$ (the part where it is deformed by $\Phi$), which also changes the background conductivity in $\widetilde{A}_{D}$ in that part of the domain via (1.1). From our results we conclude in this situation, that for
@@
In the context of Helmholtz scattering, see also the uniqueness results in [21, 6] for inclusion detection with some anisotropic coefficients.
1.1. The monotonicity method and related results
The reconstruction methods we present are based on the monotonicity of ND maps via the following operator inequalities (Proposition 5.5):
@@ 1.1. The monotonicity method and related results
where $\Lambda_{j}=\Lambda(A_{j})$ and $u_{2}$ is the electric potential for conductivity $A_{2}$ and Neumann condition $f$. The definiteness properties of $A_{2}A_{1}^{-1}(A_{2}-A_{1})$ turn out to be the same as for $A_{2}-A_{1}$ (Proposition 5.6). The main idea in the outer approach of the monotonicity method, given in our main results Theorems 3.6 and 3.7, is to construct test-operators that use the monotonicity to check if a given test-set $C$ fully contains $D^{\bullet}$ or not. This leads to an if and only if statement related to operator inequalities with the test-operators and the ND map for the unknown perturbed conductivity. In Theorem 3.7 the test-operators are linearized, which is beneficial for numerical implementation.
@@
Under much stronger assumptions, we also give an inner approach of the monotonicity method in Theorems 4.2 and  4.3, which checks if a given open set $B$ is contained in $D^{\bullet}$.
@@
Certain reconstruction methods for the partial data isotropic Calderón problem also rely on the same type of methodology. For piecewise constant layered conductivities, there is an exact reconstruction method for both conductivity coefficient and the piecewise constant partition [11, 12]. In a finite-dimensional settting (where the partition is known), a general piecewise constant conductivity is reconstructed in [13], and likewise via a reformulation to a convex semidefinite optimization problem in [23].
@@ 1.2. Remarks on notation
All considered vector spaces are complex; this includes the classes of conductivity coefficients. At all occurrences of using support/supremum/infimum, we make use of the essential support/supremum/infimum.
We use the convention that inner products on complex Hilbert spaces are linear in the first entry and anti-linear in the second. We denote the Euclidean inner product as $z_{1}\cdot\overline{z_{2}}$ for $z_{1},z_{2}\in\mathbb{C}^{d}$, in particular the “dot” is bilinear. The Euclidean norm is denoted $\lvert\,\cdot\,\rvert$.
@@ 2. The forward problem
for all $\xi\in\mathbb{C}^{d}$; there are other classes of strongly elliptic coefficients (when restricting to $\xi\in\mathbb{R}^{d}$) that we do not consider here, see e.g. [34, chapter 4]. We also have the subset where $A$ is self-adjoint, i.e. when $A=A^{\textup{R}}$:
@@ A Poincaré inequality, related to the mean-free condition in $H^{1}_{\diamond}(\Omega)$, enables that
and the forward problem is the nonlinear mapping $\Lambda\colon\mathcal{H}(\Omega)\to\mathscr{L}(L^{2}_{\diamond}(\Gamma))$. From (2.2) it holds that for any $A_{1},A_{2}\in\mathcal{H}(\Omega)$ and $f,g\in L^{2}_{\diamond}(\Gamma)$, we have
@@
$\langle f,\Lambda(A_{1})^{*}g\rangle=\overline{\langle g,\Lambda(A_{1})f\rangle}=\int_{\Omega}\overline{A_{2}\nabla u_{g}^{A_{2}}}\cdot\nabla u_{f}^{A_{1}}\,\mathrm{d}x.$
@@
Specifically,
$\Lambda(A)^{*}=\Lambda(A^{*}),$
meaning that $\Lambda(A)$ is self-adjoint if $A\in\mathcal{H}_{\textup{SA}}(\Omega)$. We also write
@@ 3. Outer reconstruction of general anisotropic inclusions
The outer shape $D^{\bullet}$ of $D$, is the smallest closed set with connected complement such that $D\subseteq D^{\bullet}$.
@@
Suppose $U\subseteq\overline{\Omega}$ is a relatively open and connected set that intersects $\Gamma$. We say that $A\in\mathcal{H}(\Omega)$ satisfies the weak unique continuation principle (UCP) in $U$ for the conductivity equation, provided that only the trivial solution of
@@
can be identically zero in a non-empty open subset of $U$, and only the trivial solution has vanishing Cauchy data, $u$ and $\nu\cdot(A\nabla u)$, on a non-empty open part of $\partial U\cap\Gamma$. If $U=\overline{\Omega}$ we just say that $A$ satisfies the UCP.
Remark $3.$
@@
We note that there also exist UCP results related to classes of complex coefficients, e.g. with principally normal symbol [31, chapter 28].
@@
    Assume there are known bounds on $A_{D}$, i.e. scalars $0<\alpha<\beta$ such that
@@
    Assume for every open neighborhood $W$ of $x\in\partial D^{\bullet}$, there exists a relatively open connected set $V\subset D^{\bullet}\cap W$ that intersects $\partial D^{\bullet}$, satisfying either of two options:
@@
    Note that Assumption 3.4(iii) only relates to what happens close to $\partial D^{\bullet}$, so further into $D$ we may have that $A_{D}-A_{0}$ is e.g. indefinite. Moreover, the definiteness condition is not exactly at $\partial D^{\bullet}$ but just close to $\partial D^{\bullet}$, meaning that there can be a continuous deviation from $A_{0}$ rather than a jump condition; in the isotropic case this mimics the condition of a strict local increase/decrease from the background conductivity when entering $D$ from the outside [15].
@@
We also introduce linearized versions of the test-operators, which are more suitable for numerical computations. Note that for $A\in\mathcal{H}_{\textup{SA}}(\Omega)$ and $B\in L^{\infty}(\Omega)^{d\times d}$, the Frechét derivative of $\Lambda$ evaluated at $A$ and in direction $B$ is given by (c.f. [14]):
@@
Pick $0<\alpha<\beta$ such that Assumption 3.4(i) holds for both $A_{D}$ and $A_{0}$. For any measurable $C\subseteq\overline{\Omega}$ we have
@@ 4. Inner reconstruction of uniformly definite anisotropic inclusions
$\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\quad\text{implies}\quad B\subseteq D^{\bullet}.$
@@
$\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0})\quad\text{implies}\quad B\subseteq D^{\bullet}.$
@@
$D^{\circ}\subseteq\cup\,\{\,B\subset\Omega\text{ open ball}\mid\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\,\}\subset D^{\bullet},$
@@
$D^{\circ}\subseteq\cup\,\{\,B\subset\Omega\text{ open ball}\mid\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0})\,\}\subset D^{\bullet}.$
@@ 5. Operator inequalities for Neumann-to-Dirichlet maps
Although we will only consider coefficients from $\mathcal{H}_{\textup{SA}}(\Omega)$ in the reconstruction methods, we still prove the following more general inequalities based on $\mathcal{H}(\Omega)$ for possible future reference. See also [24, Lemma 2.1] for the isotropic case with complex conductivity.
@@
The semidefinite case ($c=0$) of (5.7) is a simple consequence of $t\mapsto-t^{-1}$ being an operator monotone function, c.f. [8, chapter 7]. To prove (5.7), first note that we have

### v2

@@ Abstract.
We extend the monotonicity method for direct exact reconstruction of inclusions in the partial data Calderón problem, to the case of general anisotropic conductivities in any spatial dimension $d\geq 2$. From a local Neumann-to-Dirichlet map, we give reconstruction methods of inclusions based on unknown anisotropic self-adjoint perturbations to a known anisotropic conductivity coefficient. This additionally provides new insights into the non-uniqueness issues of the anisotropic Calderón problem.
The main assumption is a definiteness condition for the perturbations near the outer inclusion boundaries. Beyond this condition, they are $L^{\infty}$-perturbations that may be indefinite away from the outer inclusion boundaries, and with no boundary regularity requirement for the inclusions. Alternatively, we allow extreme parts that are perfectly insulating or perfectly conducting, in which case we require Lipschitz regularity of the outer inclusion boundaries.
@@ 1. Introduction
produces the same local Neumann-to-Dirichlet (ND) map as the coefficient $A$. For $\Gamma=\partial\Omega$ and with spatial dimension $d=2$, the obstructions to uniqueness are characterized by (1.1) for real-valued symmetric $L^{\infty}$-conductivities [4, 38]. In spatial dimension $d\geq 3$, the full extent of the non-uniqueness issues (beyond real-analytic coefficients) remains one of the most important unanswered questions in the inverse problems community [33, 40]. In a fractional Calderón problem, using the stronger non-local properties, the obstructions to uniqueness are characterized for $d\geq 2$ [10].
Despite knowledge on the possible obstructions, the non-uniqueness has prevented development of practically useful reconstruction methods. Unique recovery is only guaranteed with very strong restrictions on the considered class of coefficients [2, 3, 11].
@@
from knowledge of $A_{0}$ and the local ND map $\Lambda(A_{D})$ associated with the unknown conductivity $A_{D}$. Here $D^{\bullet}$ is the smallest closed set with connected complement such that $D\subseteq D^{\bullet}$. We give direct reconstruction methods (Theorems 3.6, 3.7, and 12.2) based on the monotonicity of the forward map $A\mapsto\Lambda(A)$. The main assumptions are self-adjointness of the coefficients, that $A_{0}$ satisfies a unique continuation principle, and that $A_{D}-A_{0}$ satisfies a local uniform definiteness condition inside $D$ near $\partial D^{\bullet}$; see Assumption 3.4 for the precise statement. In particular, inside $D$ we do not assume more regularity than $L^{\infty}$ for $A_{D}$, and $A_{D}-A_{0}$ can even be indefinite at a positive distance away from $\partial D^{\bullet}$.
As a consequence, for a given self-adjoint “background conductivity” $A_{0}$ satisfying the unique continuation principle, and any self-adjoint conductivities $A_{1}$ and $A_{2}$ (that equal $A_{0}$ near $\partial\Omega$), we have that
[…]
Hence, the obstructions from the anisotropic Calderón problem do not affect the inclusion detection problem under the definiteness condition. Practical reconstruction of inclusions is therefore sensible in a very high generality compared to reconstructing an anisotropic conductivity.
To provide some further intuition, consider a self-adjoint conductivity $A_{D}$, satisfying the definiteness condition, and with $\widetilde{A}_{D}$ defined from $A_{D}$ via (1.1) for a transformation $\Phi$. Let
@@
be the transformed inclusion under $\Phi$. If $\widetilde{D}^{\bullet}\neq D^{\bullet}$ then $\Phi$ will fail to be the identity in an open set intersecting $\partial D^{\bullet}$ (the part where it is deformed by $\Phi$), which also changes the background conductivity in $\widetilde{A}_{D}$ in that part of the domain via (1.1). We conclude for
@@
See also [8] for a stability result in a particular geometric setting. Moreover for Helmholtz scattering, there are uniqueness results for inclusion detection with some anisotropic coefficients [22, 6].
1.1. Outlining the main results
The reconstruction methods we present are based on monotonicity of the forward map via the following operator inequalities (Proposition 5.5):
@@ 1.1. Outlining the main results
where $\Lambda_{j}=\Lambda(A_{j})$ are the local ND maps, and $u_{2}$ is the electric potential for conductivity $A_{2}$ and Neumann condition $f$. The definiteness properties of $A_{2}A_{1}^{-1}(A_{2}-A_{1})$ turn out to be the same as for $A_{2}-A_{1}$ (Proposition 5.6).
The outer approach of the monotonicity method, given in our main results (Theorems 3.6, 3.7, and 12.2), constructs test-operators that use the monotonicity to check if a given admissible test-set $C$ fully contains $D^{\bullet}$ or not. Reconstruction is of the form (from Theorem 3.6):
$D\subseteq C\quad\text{\emph{if and only if}}\quad\Lambda_{C}^{-}\geq\Lambda(A_{D})\geq\Lambda_{C}^{+},$
for test-operators $\Lambda_{C}^{-}$ and $\Lambda_{C}^{+}$, and the local ND map $\Lambda(A_{D})$ for the unknown conductivity. In Theorem 3.7 the test-operators are linearized, which is beneficial for numerical implementation.
For Theorem 12.2 we also allow perfectly insulating and perfectly conducting parts in the unknown conductivity, and this is also used for the test-operators, which means that no conductivity bounds are needed for this approach. Theorem 12.2 is arguably the overall main result of the paper, although with the downside of requiring Lipschitz regularity of $\partial D^{\bullet}$ (which is unlike in Theorems 3.6 and 3.7, where no explicit boundary regularity is needed for the inclusions).
The paper is written in two parts, with Sections 210 dedicated to the finite perturbation setting, while Sections 1116 are dedicated to the extreme perturbation setting, following the approach for the isotropic case in [7, 16]. It should be noted that Theorem 12.2 cannot simply be obtained via a limit from Theorem 3.6. There is no short or straightforward proof for the difficult direction in the if and only if result, and it requires the introduction of several tools specific for extreme anisotropic coefficients; see [7, remark 5.4] for additional details in the isotropic case.
@@
Under much stronger assumptions, we also give an inner approach of the monotonicity method in Theorems 4.2, 4.3, 12.4, and 12.5, which checks if a given open set $B$ is contained in $D^{\bullet}$.
1.2. Previous results for the monotonicity method
@@
Certain reconstruction methods for the partial data isotropic Calderón problem also rely on the same type of methodology. For piecewise constant layered conductivities, there is an exact reconstruction method for both conductivity coefficient and the piecewise constant partition [12, 13]. In a finite-dimensional setting (where the partition is known), a general piecewise constant conductivity is reconstructed in [14], and likewise via a reformulation to a convex semidefinite optimization problem in [24].
@@ 1.3. Remarks on notation
All considered vector spaces are complex; this includes the classes of conductivity coefficients. We use the convention that inner products on complex Hilbert spaces are linear in the first entry and anti-linear in the second. We denote the Euclidean inner product as $z_{1}\cdot\overline{z_{2}}$ for $z_{1},z_{2}\in\mathbb{C}^{d}$, in particular the “dot” is bilinear. The Euclidean norm is denoted $\lvert\,\cdot\,\rvert$.
@@
We consistently make use of the essential support/supremum/infimum. We denote the interior of a set $C\subseteq\mathbb{R}^{d}$ as $C^{\circ}$. At certain occurrences (mainly from Section 11 and onwards) we denote by $K>0$ a generic constant that may change from line to line in the mathematical contents.
@@ 2. The forward problem
for all $\xi\in\mathbb{C}^{d}$; there are other classes of strongly elliptic coefficients (when restricting to $\xi\in\mathbb{R}^{d}$) that we do not consider here, see e.g. [35, chapter 4]. We also have the subset where $A$ is self-adjoint, i.e. when $A=A^{\textup{R}}$:
@@ A Poincaré inequality, related to the mean-free condition in $H^{1}_{\diamond}(\Omega)$, enables that
and the forward problem is the nonlinear mapping $\Lambda\colon\mathcal{H}(\Omega)\to\mathscr{L}(L^{2}_{\diamond}(\Gamma))$.
From (2.2) it holds that for any $A_{1},A_{2}\in\mathcal{H}(\Omega)$ and $f,g\in L^{2}_{\diamond}(\Gamma)$, we have
@@
$\langle f,\Lambda(A_{1})^{*}g\rangle=\overline{\langle g,\Lambda(A_{1})f\rangle}=\int_{\Omega}\overline{A_{2}\nabla u_{g}^{A_{2}}}\cdot\nabla u_{f}^{A_{1}}\,\mathrm{d}x=\int_{\Omega}A_{2}^{*}\nabla u_{f}^{A_{1}}\cdot\overline{\nabla u_{g}^{A_{2}}}\,\mathrm{d}x,$
and therefore
$\langle f,\Lambda(A^{*})g\rangle=\int_{\Omega}A\nabla u_{f}^{A}\cdot\overline{\nabla u_{g}^{A^{*}}}\,\mathrm{d}x=\langle f,\Lambda(A)^{*}g\rangle.$
Specifically $\Lambda(A^{*})=\Lambda(A)^{*}$, meaning that $\Lambda(A)$ is self-adjoint if $A\in\mathcal{H}_{\textup{SA}}(\Omega)$.
@@
We also write
@@ 3. Outer reconstruction of general anisotropic inclusions
The outer shape $D^{\bullet}$ of $D$ is the smallest closed set with connected complement such that $D\subseteq D^{\bullet}$.
@@
Let $U\subseteq\overline{\Omega}$ be relatively open and connected. $A\in\mathcal{H}(\Omega)$ satisfies the weak unique continuation principle (UCP) in $U$ for the conductivity equation if only the trivial solution of
@@
can be identically zero in a non-empty open subset of $U$, and only the trivial solution has vanishing Cauchy data, $u$ and $\nu\cdot(A\nabla u)$, on a non-empty open part of $U\cap\partial\Omega$. If $U=\overline{\Omega}$ we just say that $A$ satisfies the UCP.
Remark 3.3.
It should be noted that the “UCP from Cauchy data” follows from the “UCP from open subsets”, so for simplicity we are grouping both concepts together in Definition 3.2.
@@
We note that there also exist UCP results related to classes of complex coefficients, e.g. with principally normal symbol [32, chapter 28].
@@
    Assume there are known bounds on $A_{D}$, i.e. scalars $0<\alpha\leq\beta$ such that
@@
    For every $x\in\partial D^{\bullet}$ and every open neighborhood $W$ of $x$, assume there exists a relatively open connected set $V\subset D^{\bullet}\cap W$ that intersects $\partial D^{\bullet}$, satisfying either of two options:
@@
    Note that Assumption 3.4(iii) only relates to what happens close to $\partial D^{\bullet}$, so further into $D$ we may have that $A_{D}-A_{0}$ is e.g. indefinite. Moreover, the definiteness condition is not exactly at $\partial D^{\bullet}$ but just close to $\partial D^{\bullet}$. This allows for both a jump condition or a continuous deviation from $A_{0}$; in the isotropic case this mimics the condition of a strict local increase/decrease from the background conductivity when entering $D$ from the outside [16].
@@
We also introduce linearized versions of the test-operators, which are more suitable for numerical computations. Note that for $A\in\mathcal{H}_{\textup{SA}}(\Omega)$ and $B\in L^{\infty}(\Omega)^{d\times d}$, the Fréchet derivative of $\Lambda$ evaluated at $A$ and in direction $B$ is given by (c.f. [15]):
@@
Pick $0<\alpha\leq\beta$ such that Assumption 3.4(i) holds for both $A_{D}$ and $A_{0}$. For any measurable $C\subseteq\overline{\Omega}$ we have
@@ 4. Inner reconstruction of uniformly definite anisotropic inclusions
$\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\quad\text{implies}\quad B\subset D^{\bullet}.$
@@
$\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0})\quad\text{implies}\quad B\subset D^{\bullet}.$
@@
$D^{\circ}\subseteq\cup\,\{\,B\subseteq\Omega\text{ open ball}\mid\Lambda(A_{D})-\Lambda(A_{0})\leq\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}\,\}\subset D^{\bullet},$
@@
$D^{\circ}\subseteq\cup\,\{\,B\subseteq\Omega\text{ open ball}\mid\textup{D}\mkern-1.5mu\widehat{\Lambda}^{-}_{B}\leq\Lambda(A_{D})-\Lambda(A_{0})\,\}\subset D^{\bullet}.$
@@
Remark 4.5.
In the isotropic case, one can replace the factor $(\frac{\alpha}{\beta})^{2}$ in $\textup{D}\mkern-1.5mu\widehat{\Lambda}^{+}_{B}$ by the larger factor $\frac{\alpha}{\beta}$. This is because the same replacement holds in Proposition 5.6(ii) for isotropic coefficients, see e.g. [30, example 4.4].
@@ 5. Operator inequalities for Neumann-to-Dirichlet maps
Although we will only consider coefficients from $\mathcal{H}_{\textup{SA}}(\Omega)$ in the reconstruction methods, we still prove the following more general inequalities based on $\mathcal{H}(\Omega)$ for possible future reference. See also [25, Lemma 2.1] for the isotropic case with complex conductivity.
@@
The semidefinite case ($c=0$) of (5.7) is a simple consequence of $t\mapsto-t^{-1}$ being an operator monotone function, c.f. [9, chapter 7]. To prove (5.7), first note that we have